# Fairness in Social Influence Maximization via Optimal Transport

### v3

@@ E.1 Pseudocode
8:function s3d_step(seedset, G, p, fair_to_efficacy) $\triangleright$ each step delivers a new seedset
@@ E.2 Estimating Runtime
We estimate the running time of Algorithm 2 and 3 combined. For the S3D_STEP, lines $9$-$13$ are constant operations and comprise dataset properties. Line $14,15$ cost $O(|S|)$. FIT_TO_SIZE can cost up to $O(|S|\log|V|)$ for sampling new $|S|$ nodes. SEEDSET_REACH does repeated BFS, and so costs $O(R(|V|+|E|))$. Lines $19$-$24$ cost as follows,
$)O((|S-1)(Rd_{\text{avg}}^{D_{\text{max}}}+R|V|+\log{R|V|})$
@@
Here, we used the assumption that $d_{\text{avg}}=O(2E/V)=O(1)$ for a sparse graph ($E=O(V)$). Now this S3D_STEP is run $k$ times using S3D_ITERATE to find the best seedset in these $k$ runs. Moreover, we avoid any redundant calculations and memorize $\beta$-fariness for any seedset we discover. Hence, the total runtime is $O(kR|S||V|)$, as claimed.
@@ E.4 Theoretical Guarantees of Convergence in S3D
The S3D algorithm is similar to non-convex optimization methods such as Simulated Annealing. Such algorithms do not have theoretical guarantees but do have a long history of empirical success.
Let $f:\mathcal{P}(V)\rightarrow[0,1]$ be the $\beta$-fairness set-evaluation function defined in the power-set of the graph vertex set $V$. The function can then evaluate any seedset, $S\subseteq V$ for its $\beta$-fairness. Now for iterative optimization purposes, S3D defines a sampling process to define neighbors $S^{\prime}$ of $S$, based on similar outreaches $V_{S^{\prime}}$ and $V_{S}$. Then S3D essentially followsnon-convex optimization of $f$ using Simulated Annealing under Metropolis Sampling at a constant temperature. While Simulated Annealing does not have strict mathematical guarantees to find the global optimum in finite time, its empirical success is well understood in non-convex optimization.
While Simulated Annealing usually runs for finite iterations defined by an empirically tested temperature schedule, we ran Simulated Annealing under several constant temperatures to estimate the performance of S3D against baselines and concluded that a number of iterations $k\in[500,1000]$ usually works well in practice. Hence any decaying temperature schedule that translates total iterations in this range should work fine.
@@ E.5 Illustrative Example
Consider the information spreading over the graph in Fig. 15 as an independent cascade model with probability $p=0.1$, with blue and red nodes belonging to two different groups. A greedy strategy would choose the seed set as $S_{g}={3,5}$ (enlarged nodes) as shown in Fig. 15(a), thus leading to the highly unfair outreach in Fig. 15(b). On the contrary, our algorithm S3D promotes the choice $S_{\texttt{S3D}}={1,4}$ reflected in 15(c), which gives the more fair outreach plotted in Fig. 15(d), showing that it improves over greedy/sophisticated label-blind seed selection strategies.
@@
Figure 15: Toy example to show label-aware choice using S3D over a label-blind seedset selection process. The enlarged nodes are selected seeds. Since the graph is small, the outreach discretization bucket have been granularized for improved readability.
@@ Appendix G Declaration of Computational Resources
All experiments were performed on a local PC on a single CPU core $3.5$ GHz. Except for datasets DZ, INS, all datasets were loaded and operated on a local PC with $32$ GB of RAM. For the largest datasets (DZ, INS), we used remote compute clusters with $\sim 64$ GB memory and similar CPU capabilities. For the code development, we broadly used Python 3.10+, numpy, jupyter, and networkx [9]. Runtime for each non-S3D configured experiment on datasets except DZ, INS, was $10-15$ minutes. For DZ, INS, this was approximately $1-2$ hours. For S3D optimizations to be satisfactory, we ran each small dataset (except DZ, INS) for $1.5$ hours additionally. For massive datasets DZ, INS, the compute cluster took $\sim 4$ days for $k=10$ steps. The total set of experiments made, including the failed and passed or submitted ones, roughly took the same order of resources separately.

### v4

@@ E.1 Pseudocode
8:function s3d_step(seedset, G, p, fair_to_efficacy) $\triangleright$ each step yields a new seedset
@@ E.2 Estimating Runtime
We estimate the running time of Algorithm 2 and 3 combined. For the S3D_STEP, lines $9$-$13$ are constant operations and comprise dataset properties. Lines $14$-$15$ cost $O(|S|)$. FIT_TO_SIZE can cost up to $O(|S|\log|V|)$ for sampling new $|S|$ nodes. SEEDSET_REACH does repeated BFS, and so costs $O(R(|V|+|E|))$. Lines $19$-$24$ cost as follows,
$O((|S-1|)(Rd_{\text{avg}}^{D_{\text{max}}}+R|V|+\log{R|V|}))$ )
@@
Here, we used the assumption that $d_{\text{avg}}=O(2E/V)=O(1)$ for a sparse graph ($E=O(V)$). Now this S3D_STEP is run $k$ times using S3D_ITERATE to find the best seedset in these $k$ runs. Moreover, we avoid any redundant calculations and memorize $\beta$-fairness for any seedset we discover. Hence, the total runtime is $O(kR|S||V|)$, as claimed.
@@ E.4 Theoretical Guarantees of Convergence in S3D
The S3D algorithm is similar to non-convex optimization methods such as Simulated Annealing. Such algorithms do not have theoretical guarantees but have a long history of empirical success.
Let $f:\mathcal{P}(V)\rightarrow[0,1]$ be the $\beta$-fairness set-evaluation function defined in the power set of the graph vertex set $V$. The function can then evaluate any seedset, $S\subseteq V$ for its $\beta$-fairness. Now for iterative optimization purposes, S3D defines a sampling process to define neighbors $S^{\prime}$ of $S$, based on similar outreaches $V_{S^{\prime}}$ and $V_{S}$. Then S3D essentially follows non-convex optimization of $f$ using Simulated Annealing under Metropolis Sampling at a constant temperature. While Simulated Annealing does not have strict mathematical guarantees to find the global optimum in finite time, its empirical success is well understood in non-convex optimization.
While Simulated Annealing usually runs for finite iterations defined by an empirically tested temperature schedule, we ran Simulated Annealing under several constant temperatures to estimate the performance of S3D against baselines and concluded that a number of iterations $k\in[500,1000]$ usually works well in practice. Hence, any decaying temperature schedule that translates total iterations in this range should work fine.
@@ E.5 Illustrative Example
Consider the information spreading over the graph in Fig. 15 as an independent cascade model with probability $p=0.1$, with blue and red nodes belonging to two different groups. A greedy strategy would choose the seed set as $S_{g}={3,5}$ (enlarged nodes) as shown in Fig. 15(a), thus leading to the highly unfair outreach in Fig. 15(b). In contrast, our algorithm S3D promotes the choice $S_{\texttt{S3D}}={1,4}$ reflected in 15(c), which gives the more fair outreach plotted in Fig. 15(d), showing that it improves over greedy/sophisticated label-blind seed selection strategies.
@@
Figure 15: Toy example to show label-aware choice using S3D over a label-blind seedset selection process. The enlarged nodes are selected seeds. Since the graph is small, the outreach discretization bucket has been granularized for improved readability.
@@ Appendix G Declaration of Computational Resources
All experiments were performed on a local PC on a single CPU core $3.5$ GHz. Except for datasets DZ, INS, all datasets were loaded and operated on a local PC with $32$ GB of RAM. For the largest datasets (DZ, INS), we used remote compute clusters with $\sim 64$ GB memory and similar CPU capabilities. For the code development, we broadly used Python 3.10+, numpy, jupyter, and networkx [9]. Runtime for each non-S3D configured experiment on datasets except DZ, INS, was $10-15$ minutes. For DZ, INS, this was approximately $1-2$ hours. For S3D optimizations to be satisfactory, we ran each small dataset (except DZ, INS) for $1.5$ hours additionally. For massive datasets DZ, INS, the cluster took $\sim 4$ days for $k=10$ steps. The total set of experiments made, including the failed, passed or submitted ones, roughly took the same order of resources separately.